# Leave-One-Out Analysis for Nonconvex Robust Matrix Completion with General Thresholding Functions

### v1

@@ Appendix A Useful Lemmas and Supplementary Proofs
In this section we first collect some useful lemmas which are used throughout this paper. Those results are mostly from the literature and proofs for the new ones are provided.
@@ A.1 Useful Lemmas
Consider $m$ independent random matrices $M_{l}$ $(1\leq l\leq m)$ of dimension $d_{1}\times d_{2}$, each satisfying $\mathbb{E}\left[M_{l}\right]=0$ and $\left\|M_{l}\right\|\leq B$. Define
$\sigma^{2}:=\max\left\{\left\|\sum_{l=1}^{m}\mathbb{E}\left[M_{l}M_{l}^{T}\right]\right\|,\left\|\sum_{l=1}^{m}\mathbb{E}\left[M_{l}^{T}M_{l}\right]\right\|\right\}.$
@@
$\left\|\sum_{l=1}^{m}M_{l}\right\|\leq c_{10}\left(\sqrt{a\sigma^{2}\log\left(d_{1}+d_{2}\right)}+aB\log\left(d_{1}+d_{2}\right)\right)$
@@
Lemma 13 ([9, Lemma 4]).
Suppose that each $(i,j)$ is included in $\Omega_{0}\subseteq[n]\times[n]$ independently with probability $p_{0}$, and that $p_{0}\geq C_{13}\frac{\mu r\log n}{n}$ for some sufficiently large universal constant $C_{13}>0$. Then with high probability,
$\frac{1}{32\kappa}\|H\|_{\mathrm{F}}^{2}\leq\frac{1}{p_{0}}\left\|\mathcal{P}_{\Omega_{0}}(H)\right\|_{\mathrm{F}}^{2}\leq 40\kappa\|H\|_{\mathrm{F}}^{2},\quad\text{ for all }H\in T$
holds simultaneously for all $(X,Y)$ obeying
$\max\left\{\left\|X-U^{\star}\left(\Sigma^{\star}\right)^{1/2}\right\|_{2,\infty},\left\|Y-V^{\star}\left(\Sigma^{\star}\right)^{1/2}\right\|_{2,\infty}\right\}\leq\frac{c_{13}}{\kappa\sqrt{n}}\left(\sigma_{1}^{\star}\right)^{1/2}.$
Here, $c_{13}>0$ is some sufficiently small universal constant, and $T$ denotes the tangent space of the set of rank-r matrices at the point $XY^{T}$.

### v2

@@ Appendix A Useful Lemmas and Supplementary Proofs
In this section we collect some useful lemmas which are used throughout this paper. Proofs are mostly provided for the new ones and when the result is different from its reference.
@@ A.1 Useful Lemmas
Proof.
Denote by $\left\|\cdot\right\|_{1}$ the operator norm induced by the $l_{1}$ norm. The result follows from
$\left\|S\right\|_{2}\leq\sqrt{\left\|S\right\|_{1}\left\|S^{T}\right\|_{1}},$
and the fact that
$\max\left\{\left\|S\right\|_{1},\left\|S^{T}\right\|_{1}\right\}\leq\alpha n\cdot\left\|S\right\|_{\infty}.$
@@
Consider $m$ independent random matrices $M_{l}$ $(1\leq l\leq m)$ of dimension $d_{1}\times d_{2}$, each satisfying $\mathbb{E}\left[M_{l}\right]=0$ and $\left\|M_{l}\right\|_{2}\leq B$. Define
$\sigma^{2}:=\max\left\{\left\|\sum_{l=1}^{m}\mathbb{E}\left[M_{l}M_{l}^{T}\right]\right\|_{2},\left\|\sum_{l=1}^{m}\mathbb{E}\left[M_{l}^{T}M_{l}\right]\right\|_{2}\right\}.$
@@
$\left\|\sum_{l=1}^{m}M_{l}\right\|_{2}\leq c_{10}\left(\sqrt{a\sigma^{2}\log\left(d_{1}+d_{2}\right)}+aB\log\left(d_{1}+d_{2}\right)\right)$
@@
The following lemma is the main tool we use to bound the outlier term in the induction steps.
Lemma 13.
Recall that $\Omega_{S^{\star}}$ is the support of the outliers in $\Omega$. If $\mathcal{P}_{\Omega}(S^{\star})$ is $2\alpha p$-sparse, then
$\left\|\mathcal{P}_{\Omega_{S^{\star}}}(AB^{T})\right\|_{\mathrm{F}}^{2}\leq 2\alpha pn\cdot\min\left\{\left\|A\right\|_{\mathrm{F}}^{2}\left\|B\right\|_{2,\infty}^{2},\left\|A\right\|_{2,\infty}^{2}\left\|B\right\|_{\mathrm{F}}^{2}\right\}$
holds uniformly for all $A,B\in\mathbb{R}^{n\times r}$.
Proof.
In the following we prove that $\left\|\mathcal{P}_{\Omega_{S^{\star}}}(AB^{T})\right\|_{\mathrm{F}}^{2}\leq 2\alpha pn\cdot\left\|A\right\|_{\mathrm{F}}^{2}\left\|B\right\|_{2,\infty}^{2}$. The other part can be obtained similarly. By a direct calculation, one has
$\left\|\mathcal{P}_{\Omega_{S^{\star}}}(AB^{T})\right\|_{\mathrm{F}}^{2}=$ $\sum_{i=1}^{n}\sum_{(i,j)\in\Omega_{S^{\star}}}\left(A_{i,:}\left(B_{j,:}\right)^{T}\right)^{2}$
$\leq$ $\sum_{i=1}^{n}\sum_{(i,j)\in\Omega_{S^{\star}}}\left\|A_{i,:}\right\|_{2}^{2}\left\|B_{j,:}\right\|_{2}^{2}$
$=$ $\sum_{i=1}^{n}\left\|A_{i,:}\right\|_{2}^{2}\cdot\left(\sum_{(i,j)\in\Omega_{S^{\star}}}\left\|B_{j,:}\right\|_{2}^{2}\right)$
$\leq$ $\sum_{i=1}^{n}\left\|A_{i,:}\right\|_{2}^{2}\cdot(2\alpha pn\left\|B\right\|_{2,\infty}^{2})=2\alpha pn\cdot\left\|A\right\|_{\mathrm{F}}^{2}\left\|B\right\|_{2,\infty}^{2},$
where the second inequality follows from the sparsity assumption about $\mathcal{P}_{\Omega}(S^{\star})$. ∎